What is the missing number in this sequence?
60, 57, 51, _, 30

Answer by prateekagrawal(56)

Here the differences are.. 60-57= 3
57-51=6
and so on.
so, next difference should be 9.
51-9= 42
so, the missing number is 42.
Check: next difference must be 12.
so, 42-12= 30.
so, this is satisfying also.
Hence missing term is 42.
.
I hope this will help you. :)
.
Regards,
Prateek Agrawal